# RPRO: Ranked Preference Reinforcement Optimization for Enhancing Medical QA and Diagnostic Reasoning

### v4

@@ 1 Introduction
2 Related Work
2.1 Medical LLMs: QA, Diagnosis, and EHR
@@
2.2 Reinforcement-Learned Medical LLMs
@@
2.3 Ranking-based Optimization in Clinical Reasoning
@@ 3.2 Problem Formulation
3.3 Task-Adaptive Chain-of-Thought Generation
@@
3.4 Medical Prompt Instruction and Reasoning
Previous research shows that effective prompt template instructions can improve model performance on CoT reasoning [zhu2024llms, wei2022chain, kojima2022large]. We use structured reasoning templates based on established clinical and biomedical reasoning patterns. Table 5 presents a general medical QA template that applies a four-step reasoning process, using binary classification at each step according to the validation criteria from Section 3.3. Specifically, background knowledge retrieval gathers relevant biomedical facts and principles from medical literature. Logical connection links the retrieved knowledge to the question context. The justification step synthesizes the reasoning to support the conclusion.
@@ 3.4 Medical Prompt Instruction and Reasoning
3.5 Reasoning Refinement Method
3.5.1 Multi-Dimensional Quality Assessment Evaluation
@@
3.5.2 Probabilistic Reasoning Refinement
@@
4 Training Framework
4.1 Ranking Preference Optimization Algorithm
@@
4.2 Linear Rewards From Rank
@@
4.3 KL-Divergence Regularization for Medical Reasoning
@@
$D_{KL,t}=\sum_{v\in\mathcal{V}}\pi_{\theta}(v\mid z,\tau,c_{<t})\log\frac{\pi_{\theta}(v\mid z,\tau,c_{<t})}{\pi_{ref}(v\mid z,\tau,c_{<t})}$ (12)
@@
$\mathcal{L}_{total}=\mathcal{L}_{RPRO-CoT}+\mathcal{L}_{KL}+\mathcal{L}_{\text{Linear}}$ (13)
5 Experiments
5.1 Experimental Setting
@@ 5.4 Evaluation Metrics
Table 1: Comparison of different models on PubMedQA and MedQA-USMLE under 0-shot, 1-shot, and 5-shot settings. Best results are in bold.
@@
Table 2: Comparison of different models on the Far Eastern Memorial Hospital (FEMH) clinical dataset using BERTScore-F1 and Cosine Similarity under 0-shot, 1-shot, and 5-shot settings. Best results are in bold.
@@
Table 3: Ablation study comparing different training methods and refinement strategies on PubMedQA and MedQA-USMLE under zero-shot settings. Best results are in bold..
@@
Table 4: Reasoning results from FEMH clinical dataset
Case 1: Myocarditis (post-Moderna)
@@
Main Results. Table 1 presents the performance comparison across all evaluated models under 0-shot, 1-shot, and 5-shot settings on both datasets. Our method consistently achieves strong gains across all configurations. On PubMedQA, RPRO (Gemma 2B) attains 60.23%, 61.76%, and 62.02% accuracy, with corresponding Macro F1 scores of 51.78%, 50.95%, and 53.12% in the 0-shot, 1-shot, and 5-shot settings, respectively. Compared to the best baseline (Gemma 7B), this represents improvements of up to 3.9% in accuracy and 3.8% in Macro F1, demonstrating the robustness of our refinement mechanism across different few-shot conditions. On MedQA-USMLE, our model achieves 42.38%, 46.27%, and 51.67% accuracy, with Macro F1 scores of 30.51%, 32.68%, and 35.24%. While maintaining competitive accuracy, RPRO consistently delivers higher Macro F1 than the strongest baseline (GPT-OSS 20B), with gains of +6.75%, +2.26%, and +3.13% across the three settings, highlighting better reasoning completeness and answer quality. Overall, our 2B-parameter model matches or surpasses much larger 7B–20B models, demonstrating the effectiveness and efficiency of our reasoning refinement strategy.
@@
We evaluate the semantic similarity between the model-generated diagnosis and the true assessment using BERTScore-F1 and sentence-level Cosine Similarity under 0-shot, 1-shot, and 5-shot settings. As shown in Table 2, our proposed RPRO (Gemma 2B) achieves the best performance across all configurations, obtaining BERTScore-F1 / Cosine scores of 0.879 / 0.423, 0.884 / 0.507, and 0.891 / 0.528 in the respective settings. Compared with larger medical-tuned models such as BioMistral-7B and MedAlpaca-7B, RPRO demonstrates higher semantic coherence and diagnostic consistency, highlighting strong few-shot generalization and reliable clinical reasoning despite its compact model size.
Representative reasoning examples generated by our method on the FEMH dataset are shown in Table 4. These cases illustrate how the model integrates subjective symptoms and objective findings to produce structured diagnostic reasoning that aligns closely with clinician-provided assessments.
@@
Ablation on Training Objectives and Refinement. We begin our analysis by examining how different training objectives and refinement strategies contribute to overall performance. As shown in Table 3, standard supervised fine-tuning (SFT) with refinement provides a strong baseline, while DPO with refinement yields moderate improvements. RPRO without refinement already surpasses most baselines, demonstrating the advantage of groupwise ranking under the Bradley–Terry formulation. When refinement is added, RPRO achieves the best results across both PubMedQA and MedQA-USMLE, significantly improving both Accuracy and Macro F1. These results indicate that (1) the ranking-based preference optimization and (2) the reasoning refinement component are complementary and jointly responsible for the overall effectiveness of our method. This motivates the more fine-grained hyperparameter ablations presented in the following subsections.
Ablation on $\bm{\beta}$ Values. Fig. 2 presents the performance comparison under different $\beta$ values on the PubMedQA and MedQA-USMLE datasets. We vary $\beta\in\{0.0,0.05,0.1,0.2\}$ to control the balance between supervised alignment and preference-driven optimization. As shown in the figure, moderate $\beta$ values (around 0.1) yield the best results, achieving the highest Accuracy (60.23% on MedQA-USMLE, 42.38% on PubMedQA) and Macro F1 (55.83% and 30.51%, respectively). When $\beta$ is too small, the model underfits the preference signal; when $\beta$ is too large, over-regularization reduces reasoning diversity. This indicates that an appropriate trade-off between the base objective and preference regularization is crucial for stable optimization.
@@
Ablation on Rollout Budget $\bm{K}$. To further investigate the influence of rollout budget on reasoning quality, we vary the number of rollouts $K\in{2,4,5,8}$ during preference optimization. As illustrated in Fig. 3, increasing $K$ from 2 to 5 leads to steady improvements in both Accuracy and Macro F1 on PubMedQA and MedQA-USMLE. However, when $K$ exceeds 5, performance gains saturate and even slightly decline, indicating that excessive rollouts introduce redundant samples and reward noise, increasing computational cost without meaningful benefit. Empirically, K=5K{=}5 provides the best trade-off between performance and efficiency, consistent with prior observations in RLHF and DPO frameworks where higher rollout counts yield diminishing returns.
@@
Reasoning Results from PubMedQA and MedQA-USMLE Tasks. The reasoning examples shown in Table 5 in Appendix A.2 provide representative cases generated by our model using the proposed prompt templates. In PubMedQA, the model explicitly decomposes the question, incorporates relevant biomedical background knowledge, and connects evidence to the final answer, reflecting a literature-grounded reasoning pattern. In MedQA-USMLE, the reasoning traces follow a clinical workflow, including case summarization, identification of key findings, differential diagnosis, and treatment selection, demonstrating the model’s ability to mimic expert diagnostic procedures. Across both tasks, the generated reasoning showcases coherent multi-step inference, consistent structure, and appropriate use of domain knowledge. Together, these examples show that our approach yields structured and interpretable reasoning that is well aligned with expert biomedical and clinical decision-making, offering qualitative evidence of the model’s improved reasoning quality.
6 Conclusion and Future Work
@@ A.1 Prompt Templates
A.2 Model-Generated Reasoning Cases
The following reasoning cases, shown in Table 5, present representative outputs generated by our model using the proposed prompt templates. These cases demonstrate how our approach enables structured, interpretable, and domain-specific reasoning across both biomedical and clinical question answering tasks.
@@ Appendix C Dataset Format (JSONL Example)
Below is an illustrative JSONL example (truncated for brevity):
@@
Table 5: Reasoning results for PubMedQA and MedQA-USMLE tasks.

### v5

@@ I Introduction
II Related Work
II-A Medical LLMs: QA, Diagnosis, and EHR
@@
II-B Reinforcement-Learned Medical LLMs
@@
II-C Ranking-based Optimization in Clinical Reasoning
@@ III-B Problem Formulation
III-C Task-Adaptive Chain-of-Thought Generation
@@
III-D Medical Prompt Instruction and Reasoning
Previous research shows that effective prompt template instructions can improve model performance on CoT reasoning [zhu2024llms, wei2022chain, kojima2022large]. We use structured reasoning templates based on established clinical and biomedical reasoning patterns. Table V presents a general medical QA template that applies a four-step reasoning process, using binary classification at each step according to the validation criteria from Section III-C. Specifically, background knowledge retrieval gathers relevant biomedical facts and principles from medical literature. Logical connection links the retrieved knowledge to the question context. The justification step synthesizes the reasoning to support the conclusion.
@@ III-D Medical Prompt Instruction and Reasoning
III-E Reasoning Refinement Method
III-E1 Multi-Dimensional Quality Assessment Evaluation
@@
III-E2 Probabilistic Reasoning Refinement
@@
IV Training Framework
IV-A Ranking Preference Optimization Algorithm
@@
IV-B Linear Rewards From Rank
@@
IV-C KL-Divergence Regularization for Medical Reasoning
@@
$\mathcal{D}_{KL,t}=\sum_{v\in\mathcal{V}}\pi_{\theta}(v\mid z,\tau,c_{<t})\log\frac{\pi_{\theta}(v\mid z,\tau,c_{<t})}{\pi_{ref}(v\mid z,\tau,c_{<t})}$ (12)
@@
$\mathcal{L}_{total}=\mathcal{L}_{RPRO-CoT}+\mathcal{L}_{\text{Linear}}$ (13)
V Experiments
V-A Experimental Setting
@@ V-D Evaluation Metrics
TABLE I: Comparison of different models on PubMedQA and MedQA-USMLE under 0-shot, 1-shot, and 5-shot settings. Best results are in bold.
@@
TABLE II: Comparison of different models on the Far Eastern Memorial Hospital (FEMH) clinical dataset using BERTScore-F1 and Cosine Similarity under 0-shot, 1-shot, and 5-shot settings. Best results are in bold.
@@
TABLE III: Ablation study comparing different training methods and refinement strategies on PubMedQA and MedQA-USMLE under zero-shot settings. Best results are in bold..
@@
TABLE IV: Reasoning results from FEMH clinical dataset
[…]
@@
Main Results. Table I presents the performance comparison across all evaluated models under 0-shot, 1-shot, and 5-shot settings on both datasets. Our method consistently achieves strong gains across all configurations. On PubMedQA, RPRO (Gemma 2B) attains 60.23%, 61.76%, and 62.02% accuracy, with corresponding Macro F1 scores of 51.78%, 50.95%, and 53.12% in the 0-shot, 1-shot, and 5-shot settings, respectively. Compared to the best baseline (Gemma 7B), this represents improvements of up to 3.9% in accuracy and 3.8% in Macro F1, demonstrating the robustness of our refinement mechanism across different few-shot conditions. On MedQA-USMLE, our model achieves 42.38%, 46.27%, and 51.67% accuracy, with Macro F1 scores of 30.51%, 32.68%, and 35.24%. While maintaining competitive accuracy, RPRO consistently delivers higher Macro F1 than the strongest baseline (GPT-OSS 20B), with gains of +6.75%, +2.26%, and +3.13% across the three settings, highlighting better reasoning completeness and answer quality. Overall, our 2B-parameter model matches or surpasses much larger 7B–20B models, demonstrating the effectiveness and efficiency of our reasoning refinement strategy.
@@
We evaluate the semantic similarity between the model-generated diagnosis and the true assessment using BERTScore-F1 and sentence-level Cosine Similarity under 0-shot, 1-shot, and 5-shot settings. As shown in Table II, our proposed RPRO (Gemma 2B) achieves the best performance across all configurations, obtaining BERTScore-F1 / Cosine scores of 0.879 / 0.423, 0.884 / 0.507, and 0.891 / 0.528 in the respective settings. Compared with larger medical-tuned models such as BioMistral-7B and MedAlpaca-7B, RPRO demonstrates higher semantic coherence and diagnostic consistency, highlighting strong few-shot generalization and reliable clinical reasoning despite its compact model size.
Representative reasoning examples generated by our method on the FEMH dataset are shown in Table IV. These cases illustrate how the model integrates subjective symptoms and objective findings to produce structured diagnostic reasoning that aligns closely with clinician-provided assessments.
@@
Ablation on Training Objectives and Refinement. We begin our analysis by examining how different training objectives and refinement strategies contribute to overall performance. As shown in Table III, standard supervised fine-tuning (SFT) with refinement provides a strong baseline, while DPO with refinement yields moderate improvements. RPRO without refinement already surpasses most baselines, demonstrating the advantage of groupwise ranking under the Bradley–Terry formulation. When refinement is added, RPRO achieves the best results across both PubMedQA and MedQA-USMLE, significantly improving both Accuracy and Macro F1. These results indicate that (1) the ranking-based preference optimization and (2) the reasoning refinement component are complementary and jointly responsible for the overall effectiveness of our method. This motivates the more fine-grained hyperparameter ablations presented in the following subsections.
Ablation on $\boldsymbol{\beta}$ Values. Fig. 2 presents the performance comparison under different $\beta$ values on the PubMedQA and MedQA-USMLE datasets. We vary $\beta\in\{0.0,0.05,0.1,0.2\}$ to control the balance between supervised alignment and preference-driven optimization. As shown in the figure, moderate $\beta$ values (around 0.1) yield the best results, achieving the highest Accuracy (60.23% on MedQA-USMLE, 42.38% on PubMedQA) and Macro F1 (55.83% and 30.51%, respectively). When $\beta$ is too small, the model underfits the preference signal; when $\beta$ is too large, over-regularization reduces reasoning diversity. This indicates that an appropriate trade-off between the base objective and preference regularization is crucial for stable optimization.
@@
Ablation on Rollout Budget $\boldsymbol{K}$. To further investigate the influence of rollout budget on reasoning quality, we vary the number of rollouts $K\in{2,4,5,8}$ during preference optimization. As illustrated in Fig. 3, increasing $K$ from 2 to 5 leads to steady improvements in both Accuracy and Macro F1 on PubMedQA and MedQA-USMLE. However, when $K$ exceeds 5, performance gains saturate and even slightly decline, indicating that excessive rollouts introduce redundant samples and reward noise, increasing computational cost without meaningful benefit. Empirically, K=5K{=}5 provides the best trade-off between performance and efficiency, consistent with prior observations in RLHF and DPO frameworks where higher rollout counts yield diminishing returns.
@@
Reasoning Results from PubMedQA and MedQA-USMLE Tasks. The reasoning examples shown in Table V in Appendix A-B provide representative cases generated by our model using the proposed prompt templates. In PubMedQA, the model explicitly decomposes the question, incorporates relevant biomedical background knowledge, and connects evidence to the final answer, reflecting a literature-grounded reasoning pattern. In MedQA-USMLE, the reasoning traces follow a clinical workflow, including case summarization, identification of key findings, differential diagnosis, and treatment selection, demonstrating the model’s ability to mimic expert diagnostic procedures. Across both tasks, the generated reasoning showcases coherent multi-step inference, consistent structure, and appropriate use of domain knowledge. Together, these examples show that our approach yields structured and interpretable reasoning that is well aligned with expert biomedical and clinical decision-making, offering qualitative evidence of the model’s improved reasoning quality.
VI Conclusion and Future Work
@@ A-A Prompt Templates
A-B Model-Generated Reasoning Cases
The following reasoning cases, shown in Table V, present representative outputs generated by our model using the proposed prompt templates. These cases demonstrate how our approach enables structured, interpretable, and domain-specific reasoning across both biomedical and clinical question answering tasks.
@@ Appendix C Dataset Format (JSONL Example)
TABLE V: Reasoning results for PubMedQA and MedQA-USMLE tasks.